OpenStudy (anonymous):

x varies directly with x.If x = 2 and y = 8,find the constant of variation.(show and explain the solution)

OpenStudy (anonymous):

is ur question properly put????

OpenStudy (anonymous):

sorry the question should be x varies directly with y.If x = 2 and y = 8,find the constant of variation.(show and explain the solution)

OpenStudy (anonymous):

y=k*x input data x=2 and y=8 8=k*2 k=4 so y=4x

OpenStudy (anonymous):

there is not any k in the question, how u got the answer ?????

OpenStudy (kira_yamato):

We know that x varies directly with y. y = k x, where k is just some constant. Plug in x = 2 and y = 8 8 = 2 k k = 4. Hence y = 4x (QED)

OpenStudy (anonymous):

thanksss